OpenStudy (anonymous):

the integral from 0 to 25 of dx/ square root (4+ square root x)

OpenStudy (anonymous):

\[\int\limits_{0}^{25} dx / \sqrt{4+\sqrt{x}}\]

OpenStudy (anonymous):

uses substitution respect to u

OpenStudy (anonymous):

u= \[\sqrt{4+\sqrt x}\]

OpenStudy (anonymous):

use the derivative

OpenStudy (anonymous):

then du = \[1/2(4+\sqrt{x})^{-1/2}\] dx confused on what to do then

OpenStudy (anonymous):

would it be easier to make two separate u subs?

OpenStudy (anonymous):

chain rule

OpenStudy (anonymous):

1/2sqrt(4+sqrt x )* 1/2 sqrt x

OpenStudy (anonymous):

where would you go from there?

OpenStudy (anonymous):

correct me if i am wrong, but wouldnt it be to the power of -1/2 rather than 1/2?

OpenStudy (anonymous):